# Plug-and-Hide: Provable and Adjustable Diffusion Generative Steganography

### 03 Mar 2026 v4

@@ 4.3 Adjustable PA-B2G
According to Theorem˜1,
$p(g)=\sum\limits_{m}p(g|\mathcal{C}_{u}(m))p(\mathcal{C}_{u}(m))p(m)$ (10)
@@
According to Equation˜11 and Equation˜12, the following conclusions hold true for both modes I and II: i) $gp(g)$ is an odd function; ii) the integral interval is symmetric about the axis $g=0$; and iii) the functions $gp(g)$ is bounded and contains a finite number of jump discontinuities within the integral interval. Collecting these conditions, we can derive that $\mathrm{E}[g]=\int_{-\infty}^{\infty}gp(g)\mathrm{d}g=0$. According to the Law of Large Numbers, when $k\to\infty$, we have $E[g]=\bar{g}=0$. Moreover, $D[g]=E[g^{2}]-(E[g])^{2}=E[g^{2}]=\int_{-\infty}^{+\infty}g^{2}p(g)\mathrm{d}g$. Clearly, according to Equation˜11 and Equation˜12, $D(g)=E[g^{2}]\neq 1$, and therefore, when $k\to\infty$, $D(g)=S^{2}_{g}\neq 1$. ∎
Theorem˜2 indicates that the introduction of no-sampling intervals theoretically does not change the sample mean $\bar{g}$ of $\mathbf{g}$, primarily due to the symmetry of our interval partitioning strategy proposed in Section˜4.2. However, it induces a change in the sample variance $S_{g}^{2}$. Therefore, we only need to correct $S^{2}_{g}$ to satisfy the three conditions mentioned above. To this end, we design a variance-preserving algorithm, whose essence lies in iteratively adjusting the sampling intervals of $u$ to ensure that $S^{2}_{g}$ converges to 1. Specifically, if $S^{2}_{g}>1$, we iteratively reduce the sample variance $S^{2}_{u}$ of $\mathbf{u}$; otherwise, we iteratively increase $S^{2}_{u}$. Figure˜1 shows two examples for $l=2$ w.r.t. modes I and II, respectively, illustrating how $S^{2}_{u}$ is increased or decreased during each iteration. The complete implementation of our adjustable PA-B2G is given in Algorithm˜3, and its reverse process is identical to that in Algorithm˜2. Combining Algorithm˜3 and Figure˜1, we observe that MN, MB, MC, and Multi-bits [8] are in fact four special cases of our PA-B2G.
@@
Input: Secret message $\mathbf{m}$, maximum iteration steps $N_{max}$, and error tolerance $e$.
Output: Sampling result $\mathbf{g}_{s}$.
1:Initialize $c_{1}=c_{2}=0$;
2:Generate a seed sequence $s_{1}\cdots s_{k}$.
3:for $i=1$ to $k$ do
4:  Sample $r_{i}\sim\mathcal{U}(0,1)$ with seed $s_{i}$;
@@
6:  $h_{1}=\frac{m_{i}+1}{2^{l}}-\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{m_{i}+1}{2^{l}})-\Delta_{g})$;
7:  $h_{2}=\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{m_{i}}{2^{l}})+\Delta_{g})-\frac{m_{i}}{2^{l}}$;
8:  $u_{i}=(\frac{1}{2^{l}}-h_{1}-h_{2}-c_{1}-c_{2})r_{i}+\frac{m_{i}}{2^{l}}+h_{2}$;
9:  $u_{i}=u_{i}+c_{1}$ if $\frac{m_{i}}{2}<\frac{1}{2}$ else $u_{i}+c_{2}$;
@@
11:  $h_{1}=\frac{(m_{i}+1)\ \mathrm{mod}\ 2^{l}}{2^{l+1}}-\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{(m_{i}+1)\ \mathrm{mod}\ 2^{l}}{2^{l+1}})-\Delta_{g})$;
12:  $h_{2}=\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{m_{i}}{2^{l+1}})+\Delta_{g})-\frac{m_{i}}{2^{l+1}}$;
13:  $u_{i}=(\frac{1}{2^{l+1}}-h_{1}-h_{2}-c_{1}-c_{2})r_{i}+\frac{m_{i}}{2^{l+1}}+h_{2}+c_{1}$ or $u_{i}=(\frac{1}{2^{l+1}}-h_{1}-h_{2}-c_{1}-c_{2})r_{i}+1-\frac{m_{i}+1}{2^{l+1}}+h_{2}+c_{2}$;
14:for $step=1$ to $N_{max}$ do
15:  $\mathbf{g}=\mathcal{C}_{g}(\mathbf{u})$ and calculate $S_{g}^{2}$;
16:  if 1-$S_{g}^{2}>e$ then
17:   $c_{1}=c_{1}+\Delta_{c}$;
18:  else if $S^{2}_{g}-1>e$ then
19:   $c_{2}=c_{2}+\Delta_{c}$;
20:  else break;
21:  Repeat lines 3-13;
22:Return $\mathbf{g}_{s}$.
@@
In Algorithm˜3, we do not treat $h$ as a hyperparameter; instead, we utilize $\Delta_{g}$ to dynamically determine the neighborhoods (referred to as no-sampling intervals) around each quantile. The rationale behind this is as follows: If we model the quantile as a continuous variable $q$, first, as $q\to 0$ or $1$, $\mathcal{C}_{g}(q+h)-\mathcal{C}_{g}(q-h)$ gradually increases, which significantly disrupts the Gaussianity of $\mathbf{g}$. Second, when $g=\mathcal{C}_{g}(u)$ undergoes a small numerical perturbation $\Delta_{g}$, we observe that $\Delta_{u}\approx\frac{\mathbf{d}\mathcal{C}^{-1}_{g}}{\mathbf{d}g}\Delta_{g}$. Since $\mathcal{C}_{g}$ is the PPF of $\mathcal{N}(0,1)$, $\frac{\mathbf{d}\mathcal{C}^{-1}_{g}}{\mathbf{d}g}$ is a symmetric function that first increases and then decreases over its domain. Consequently, as $q\to 0$ or $1$, $\frac{\mathbf{d}\mathcal{C}^{-1}_{g}}{\mathbf{d}g}\to 0$, leading to $\Delta_{u}\to 0$. This implies that overly long no-sampling intervals are unnecessary for $q\to 0$ or $1$. Based on this insight, we directly introduce a hyperparameter $\Delta_{g}$ to adjust the deviation of $\mathbf{g}$ from $\mathcal{N}(\mathbf{0},\mathbf{I})$, thereby enabling adaptive control over the length of the no-sampling intervals. As shown in fig.˜1 and Theorem˜2, for a given $\Delta_{g}$, as $q\to 0$ or $1$, both $h_{1}$ and $h_{2}$ approach 0, which aligns seamlessly with our theoretical perspective.
4.4 Message Hiding and Extraction with PF-ODEs
@@
After obtaining the noise $\mathbf{g}_{s}$ via PA-B2G, $\mathbf{g}_{s}$ can be reparameterized to accommodate selected diffusion models. The process of generating stego images from $\mathbf{g}_{s}$ essentially amounts to a numerical integration problem for ODEs. A variety of ODE solvers have been applied to diffusion models. In this paper, we primarily adopt the 2nd-order Heun solver [7] for stego image generation and initial noise retrieval for the following reasons: i) Universality: [7] provides a generalized ODE formulation that is compatible with most existing diffusion models using different noise schedulers (e.g., DDIM [21], VP-SDE [22], and VE-SDE [22]), without requiring additional training or fine-tuning. ii) Efficiency: Higher-order ODE solvers typically reduce local errors but also incur increased computational costs. 2nd-order Heun solver can achieve a favorable computational trade-off.
@@
However, in practice, directly adapting the 2nd-order Heun solver to diffusion models defined over discretized timesteps may lead to numerical calculation errors. For instance, when adopting the noise scheduler of DDPM, we have $\sigma(t)=t=\sqrt{1-\bar{\alpha}_{t}}$, where $\bar{\alpha}_{t}$ denotes a transformed noise scheduler defined in [4]. This gives rise to a critical issue: when attempting to retrieve $\mathbf{g}$ from $\mathbf{x}_{0}$, the Heun sampler fails because the initial noise parameter $\sigma(0)=\sqrt{1-\bar{\alpha}_{0}}=0$ (see the deterministic sampler defined in [7] for details). To resolve this problem, we propose retrieving $\mathbf{g}_{s}$ from $\mathbf{x}_{\epsilon}$ instead of $\mathbf{x}_{0}$, where $\epsilon$ is a small constant—for example, $\epsilon=10^{-6}$ as employed in this paper. This modification does not affect the generation of stego images but effectively enhances the stability of message extraction.
5 Experiments
5.1 Experimental setup
@@
  • Steganographic security is measured by detection accuracy ($Acc_{s}$) using the state-of-the-art steganalyzer UCNet [24], with 5,000 cover-stego image pairs for training and 1,000 pairs for testing. Note that the cover images are also generated by generative models but do not contain any hidden secret messages.
@@
Figure 2: Averaged histograms of noise that rejecting $H_{0}$ w.r.t. different payloads and noise sizes under $\Delta_{g}=0.02$.
3 bpp 6 bpp 9 bpp
Noise Size($\times 3$) Mode I Mode II Mode I Mode II Mode I Mode II
$\Delta_{g}\!=\!0.00$ $32\times 32$ 0.957 0.952 0.959 0.954 0.952 0.951
$64\times 64$ 0.949 0.947 0.951 0.952 0.948 0.950
$128\times 128$ 0.947 0.949 0.947 0.953 0.950 0.945
$256\times 256$ 0.946 0.950 0.949 0.951 0.951 0.948
$\Delta_{g}\!=\!0.02$ $32\times 32$ 0.849 0.872 0.835 0.844 0.845 0.795
$64\times 64$ 0.405 0.601 0.431 0.401 0.393 0.333
$128\times 128$ 0.00 0.00 0.00 0.00 0.00 0.00
$256\times 256$ 0.00 0.00 0.00 0.00 0.00 0.00
Table 1: Normality test of samples generated by PA-B2G in modes I and $I$ with four different sample sizes w.r.t. varying payloads and $\Delta_{g}$.
5.2 Normality Test
We employ the one-sample Kolmogorov-Smirnov (K-S) test to assess the Gaussianity of samples generated by PA-B2G. For each sample, we test the null hypothesis $H_{0}$, which posits that the sample follows $\mathcal{N}(0,1)$, against the alternative hypothesis $H_{1}$. A $p$-value below $0.05$ leads to the rejection of $H_{0}$; otherwise, it is accepted. Through a random generation process, we create $1,000$ samples and compute the ratio $N_{H_{0}}/1,000$, where $N_{H_{0}}$ denotes the number of samples for which $H_{0}$ is accepted. Table˜1 shows the test results w.r.t. varying $\Delta_{g}\in\{0,0.02\}$ and payloads of $3$ bpp, $6$ bpp, and $9$ bpp for both modes I and II.
As observed, when no-sampling intervals are not introduced (i.e., $\Delta_{g}=0$), the samples generated by PA-B2G for any noise size and payload almost strictly adhere to $\mathcal{N}(0,1)$. However, even though $\Delta_{g}$ is a small value, the ratio $N_{H_{0}}/1,000$ decreases significantly, particularly for large noise sizes. This may occur because the normality of samples generated by PA-B2G is highly sensitive to $\Delta_{g}$, while, conversely, the sensitivity of the K-S test improves as the sample scale increases. Furthermore, we depict the averaged histograms of noise samples that reject $H_{0}$ at varying payloads and noise sizes for $\Delta_{g}=0.02$ in Figure˜2. As illustrated, as the payload increases, the distribution of $\mathbf{g}_{s}$ gradually deviates from $\mathcal{N}(0,1)$. However, for low payloads, even when $H_{0}$ is rejected, these histograms still closely resemble the distribution of noise sampled from $\mathcal{N}(0,1)$.
5.3 Method Comparison
Method CIFAR-10 $32\times 32$ FFHQ $64\times 64$ LSUN-Bedroom $256\times 256$
3 bpp 6bpp 3 bpp 6 bpp 3 bpp 6 bpp
@@ 5.3 Method Comparison
MN [8] 2.02 83.86% 50.60% - - - 2.46 86.22% 50.83% - - - 10.21 79.46% 49.95% - - -
MB [8] 2.64 94.36% 92.35% - - - 4.16 98.12% 99.70% - - - 10.98 96.89% 99.90% - - -
MC [8] 2.25 84.96% 67.00% - - - 3.10 92.70% 97.15% - - - 10.82 82.39% 99.20% - - -
Multi-bits [8] - - - 2.37 83.01% 74.80% - - - 3.30 90.57% 98.85% - - - 10.90 80.90% 99.90%
GSD [33] 2.98 98.89% 64.54% - - - 3.08 98.81% 50.24% - - - 10.89 94.11% 50.56% - - -
GS [29] 1.99 84.50% 50.94% 2.01 74.98% 50.12% 2.42 86.24% 50.35% 2.45 79.98% 50.54% 10.21 79.84% 50.74% 10.22 72.04% 50.15%
PA-B2G1(.00) 1.97 84.10% 50.49% 2.03 75.60% 50.75% 2.45 86.85% 50.58% 2.46 79.25% 50.65% 10.19 79.46% 49.74% 10.20 71.12% 49.50%
PA-B2G1(.08) 2.19 88.16% 56.30% 2.17 80.02% 56.82% 2.95 92.76% 79.95% 2.98 88.40% 79.85% 10.75 86.10% 72.75% 10.78 78.10% 72.35%
PA-B2G2(.00) 2.01 75.64% 50.75% 2.00 69.11% 50.43% 2.40 79.22% 50.43% 2.43 70.90% 50.55% 10.21 72.14% 48.80% 10.21 62.54% 49.80%
PA-B2G2(.08) 2.13 80.99% 53.80% 2.08 71.53% 53.40% 2.94 88.35% 79.50% 2.89 77.16% 79.60% 10.71 82.87% 71.80% 10.73 71.90% 72.30%

### 04 Mar 2026 v5

@@ 4.3. Adjustable PA-B2G
According to Theorem 4.1,
(10) $p(g)=\sum\limits_{m}p(g|\mathcal{C}_{u}(m))p(\mathcal{C}_{u}(m))p(m)$
@@
According to Equation 11 and Equation 12, the following conclusions hold true for both modes I and II: i) $gp(g)$ is an odd function; ii) the integral interval is symmetric about the axis $g=0$; and iii) the functions $gp(g)$ is bounded and contains a finite number of jump discontinuities within the integral interval. Collecting these conditions, we can derive that $\mathrm{E}[g]=\int_{-\infty}^{\infty}gp(g)\mathrm{d}g=0$. According to the Law of Large Numbers, when $k\to\infty$, we have $E[g]=\bar{g}=0$. Moreover, $D[g]=E[g^{2}]-(E[g])^{2}=E[g^{2}]=\int_{-\infty}^{+\infty}g^{2}p(g)\mathrm{d}g$. Clearly, according to Equation 11 and Equation 12, $D(g)=E[g^{2}]\neq 1$, and therefore, when $k\to\infty$, $D(g)=S^{2}_{g}\neq 1$. ∎
Theorem 4.2 indicates that the introduction of no-sampling intervals theoretically does not change the sample mean $\bar{g}$ of $\mathbf{g}$, primarily due to the symmetry of our interval partitioning strategy proposed in Section 4.2. However, it induces a change in the sample variance $S_{g}^{2}$. Therefore, we only need to correct $S^{2}_{g}$ to satisfy the three conditions mentioned above. To this end, we design a variance-preserving algorithm, whose essence lies in iteratively adjusting the sampling intervals of $u$ to ensure that $S^{2}_{g}$ converges to 1. Specifically, if $S^{2}_{g}>1$, we iteratively reduce the sample variance $S^{2}_{u}$ of $\mathbf{u}$; otherwise, we iteratively increase $S^{2}_{u}$. Figure 1 shows two examples for $l=2$ w.r.t. modes I and II, respectively, illustrating how $S^{2}_{u}$ is increased or decreased during each iteration. The complete implementation of our adjustable PA-B2G is given in Algorithm 3, and its reverse process is identical to that in Algorithm 2. Combining Algorithm 3 and Figure 1, we observe that MN, MB, MC, and Multi-bits (Kim et al., 2025) are in fact four special cases of our PA-B2G.
@@
1:Secret message $\mathbf{m}$, maximum iteration steps $N_{max}$, and error tolerance $e$.
2:Sampling result $\mathbf{g}_{s}$.
3:Initialize $c_{1}=c_{2}=0$;
4:Generate a seed sequence $s_{1}\cdots s_{k}$.
5:for $i=1$ to $k$ do
6:  Sample $r_{i}\sim\mathcal{U}(0,1)$ with seed $s_{i}$;
@@
8:  $h_{1}=\frac{m_{i}+1}{2^{l}}-\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{m_{i}+1}{2^{l}})-\Delta_{g})$;
9:  $h_{2}=\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{m_{i}}{2^{l}})+\Delta_{g})-\frac{m_{i}}{2^{l}}$;
10:  $u_{i}=(\frac{1}{2^{l}}-h_{1}-h_{2}-c_{1}-c_{2})r_{i}+\frac{m_{i}}{2^{l}}+h_{2}$;
11:  $u_{i}=u_{i}+c_{1}$ if $\frac{m_{i}}{2}<\frac{1}{2}$ else $u_{i}+c_{2}$;
@@
13:  $h_{1}=\frac{(m_{i}+1)\ \mathrm{mod}\ 2^{l}}{2^{l+1}}-\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{(m_{i}+1)\ \mathrm{mod}\ 2^{l}}{2^{l+1}})-\Delta_{g})$;
14:  $h_{2}=\mathcal{C}^{-1}_{g}(\mathcal{C}_{g}(\frac{m_{i}}{2^{l+1}})+\Delta_{g})-\frac{m_{i}}{2^{l+1}}$;
15:  $u_{i}=(\frac{1}{2^{l+1}}-h_{1}-h_{2}-c_{1}-c_{2})r_{i}+\frac{m_{i}}{2^{l+1}}+h_{2}+c_{1}$ or $u_{i}=(\frac{1}{2^{l+1}}-h_{1}-h_{2}-c_{1}-c_{2})r_{i}+1-\frac{m_{i}+1}{2^{l+1}}+h_{2}+c_{2}$;
16:for $step=1$ to $N_{max}$ do
17:  $\mathbf{g}=\mathcal{C}_{g}(\mathbf{u})$ and calculate $S_{g}^{2}$;
18:  if 1-$S_{g}^{2}>e$ then
19:   $c_{1}=c_{1}+\Delta_{c}$;
20:  else if $S^{2}_{g}-1>e$ then
21:   $c_{2}=c_{2}+\Delta_{c}$;
22:  else break;
23:  Repeat lines 3-13;
24:Return $\mathbf{g}_{s}$.
@@
In Algorithm 3, we do not treat $h$ as a hyperparameter; instead, we utilize $\Delta_{g}$ to dynamically determine the neighborhoods (referred to as no-sampling intervals) around each quantile. The rationale behind this is as follows: If we model the quantile as a continuous variable $q$, first, as $q\to 0$ or $1$, $\mathcal{C}_{g}(q+h)-\mathcal{C}_{g}(q-h)$ gradually increases, which significantly disrupts the Gaussianity of $\mathbf{g}$. Second, when $g=\mathcal{C}_{g}(u)$ undergoes a small numerical perturbation $\Delta_{g}$, we observe that $\Delta_{u}\approx\frac{\mathbf{d}\mathcal{C}^{-1}_{g}}{\mathbf{d}g}\Delta_{g}$. Since $\mathcal{C}_{g}$ is the PPF of $\mathcal{N}(0,1)$, $\frac{\mathbf{d}\mathcal{C}^{-1}_{g}}{\mathbf{d}g}$ is a symmetric function that first increases and then decreases over its domain. Consequently, as $q\to 0$ or $1$, $\frac{\mathbf{d}\mathcal{C}^{-1}_{g}}{\mathbf{d}g}\to 0$, leading to $\Delta_{u}\to 0$. This implies that overly long no-sampling intervals are unnecessary for $q\to 0$ or $1$. Based on this insight, we directly introduce a hyperparameter $\Delta_{g}$ to adjust the deviation of $\mathbf{g}$ from $\mathcal{N}(\mathbf{0},\mathbf{I})$, thereby enabling adaptive control over the length of the no-sampling intervals. As shown in fig. 1 and Theorem 4.2, for a given $\Delta_{g}$, as $q\to 0$ or $1$, both $h_{1}$ and $h_{2}$ approach 0, which aligns seamlessly with our theoretical perspective.
4.4. Message Hiding and Extraction with PF-ODEs
@@
After obtaining the noise $\mathbf{g}_{s}$ via PA-B2G, $\mathbf{g}_{s}$ can be reparameterized to accommodate selected diffusion models. The process of generating stego images from $\mathbf{g}_{s}$ essentially amounts to a numerical integration problem for ODEs. A variety of ODE solvers have been applied to diffusion models. In this paper, we primarily adopt the 2nd-order Heun solver (Karras et al., 2022) for stego image generation and initial noise retrieval for the following reasons: i) Universality: (Karras et al., 2022) provides a generalized ODE formulation that is compatible with most existing diffusion models using different noise schedulers (e.g., DDIM (Song et al., 2020a), VP-SDE (Song et al., 2020b), and VE-SDE (Song et al., 2020b)), without requiring additional training or fine-tuning. ii) Efficiency: Higher-order ODE solvers typically reduce local errors but also incur increased computational costs. 2nd-order Heun solver can achieve a favorable computational trade-off.
@@
However, in practice, directly adapting the 2nd-order Heun solver to diffusion models defined over discretized timesteps may lead to numerical calculation errors. For instance, when adopting the noise scheduler of DDPM, we have $\sigma(t)=t=\sqrt{1-\bar{\alpha}_{t}}$, where $\bar{\alpha}_{t}$ denotes a transformed noise scheduler defined in (Ho et al., 2020). This gives rise to a critical issue: when attempting to retrieve $\mathbf{g}$ from $\mathbf{x}_{0}$, the Heun sampler fails because the initial noise parameter $\sigma(0)=\sqrt{1-\bar{\alpha}_{0}}=0$ (see the deterministic sampler defined in (Karras et al., 2022) for details). To resolve this problem, we propose retrieving $\mathbf{g}_{s}$ from $\mathbf{x}_{\epsilon}$ instead of $\mathbf{x}_{0}$, where $\epsilon$ is a small constant—for example, $\epsilon=10^{-6}$ as employed in this paper. This modification does not affect the generation of stego images but effectively enhances the stability of message extraction.
5. Experiments
5.1. Experimental setup
@@
  • Steganographic security is measured by detection accuracy ($Acc_{s}$) using the state-of-the-art steganalyzer UCNet (Wei et al., 2022a), with 5,000 cover-stego image pairs for training and 1,000 pairs for testing. Note that the cover images are also generated by generative models but do not contain any hidden secret messages.
@@
Figure 2. Averaged histograms of noise that rejecting $H_{0}$ w.r.t. different payloads and noise sizes under $\Delta_{g}=0.02$.
@@
Normality test of samples generated by PA-B2G in modes I and II with four different sample sizes w.r.t. varying payloads and $\Delta_{g}$. 3 bpp 6 bpp 9 bpp Noise Size($\times 3$) Mode I Mode II Mode I Mode II Mode I Mode II $\Delta_{g}\!=\!0.00$ $32\times 32$ 0.957 0.952 0.959 0.954 0.952 0.951 $64\times 64$ 0.949 0.947 0.951 0.952 0.948 0.950 $128\times 128$ 0.947 0.949 0.947 0.953 0.950 0.945 $256\times 256$ 0.946 0.950 0.949 0.951 0.951 0.948 $\Delta_{g}\!=\!0.02$ $32\times 32$ 0.849 0.872 0.835 0.844 0.845 0.795 $64\times 64$ 0.405 0.601 0.431 0.401 0.393 0.333 $128\times 128$ 0.00 0.00 0.00 0.00 0.00 0.00 $256\times 256$ 0.00 0.00 0.00 0.00 0.00 0.00
5.2. Normality Test
We employ the one-sample Kolmogorov-Smirnov (K-S) test to assess the Gaussianity of samples generated by PA-B2G. For each sample, we test the null hypothesis $H_{0}$, which posits that the sample follows $\mathcal{N}(0,1)$, against the alternative hypothesis $H_{1}$. A $p$-value below $0.05$ leads to the rejection of $H_{0}$; otherwise, it is accepted. Through a random generation process, we create $1,000$ samples and compute the ratio $N_{H_{0}}/1,000$, where $N_{H_{0}}$ denotes the number of samples for which $H_{0}$ is accepted. Section 5.1 shows the test results w.r.t. varying $\Delta_{g}\in\{0,0.02\}$ and payloads of $3$ bpp, $6$ bpp, and $9$ bpp for both modes I and II.
As observed, when no-sampling intervals are not introduced (i.e., $\Delta_{g}=0$), the samples generated by PA-B2G for any noise size and payload almost strictly adhere to $\mathcal{N}(0,1)$. However, even though $\Delta_{g}$ is a small value, the ratio $N_{H_{0}}/1,000$ decreases significantly, particularly for large noise sizes. This may occur because the normality of samples generated by PA-B2G is highly sensitive to $\Delta_{g}$, while, conversely, the sensitivity of the K-S test improves as the sample scale increases. Furthermore, we depict the averaged histograms of noise samples that reject $H_{0}$ at varying payloads and noise sizes for $\Delta_{g}=0.02$ in Figure 2. As illustrated, as the payload increases, the distribution of $\mathbf{g}_{s}$ gradually deviates from $\mathcal{N}(0,1)$. However, for low payloads, even when $H_{0}$ is rejected, these histograms still closely resemble the distribution of noise sampled from $\mathcal{N}(0,1)$.
5.3. Method Comparison